OpenStudy (anonymous):

What is the standard equation of a circle at the origin? A. x2 - y2 = r2 B. x2 + y2 = r2 C. x2 + y2 = r D. x2 - y2 = r

OpenStudy (anonymous):

B

OpenStudy (anonymous):

how do you figure that ?

OpenStudy (anonymous):

usually in a circle equation (x-h)^2+(y-k)^2= r^2 r is radius and h and k are usually points in your case origon is (0,0) leaving you with x^2+y^2=r^2